an investor will invest a total of $12000 on two accounts, one paying 5% annual simple interest and the other 3%. if she wants to earn $500 annual interest how much should she invest at each account?

Answer by lwsshak3(11628)

an investor will invest a total of $12000 on two accounts, one paying 5% annual simple interest and the other 3%. if she wants to earn $500 annual interest how much should she invest at each account?
***
x=amount invested at 5%
12000-x=amount invested at 3%
5%x+3%(12000-x)=500
.05x+360-.03x=500
.02x=140
x=7000
12000-x=5000
..
amount invested at 5%=$7000
amount invested at 3%=$5000